OpenStudy (anonymous):

solve radical 3sq.6x-9=3

myininaya (myininaya):

so we have (6x-9)^(1/3)=3 correct?

myininaya (myininaya):

to get rid of 1/3 power raise both sides to the 3rd power

myininaya (myininaya):

6x-9=3^3

OpenStudy (anonymous):

no 3cubed 6x-9=3

myininaya (myininaya):

?

myininaya (myininaya):

use that equation thing below this box

OpenStudy (anonymous):

3\sqrt{6x-9}=3\]

myininaya (myininaya):

\[\sqrt{(6x-9)^3}=3\]

myininaya (myininaya):

is that right?

OpenStudy (anonymous):

3 is index instead of 3rd power

myininaya (myininaya):

\[\sqrt[3]{6x-9}=3\]

OpenStudy (anonymous):

\[3\sqrt{6x-9}=3\]

myininaya (myininaya):

oh 3 times that okay

OpenStudy (anonymous):

sorry I am new at this

myininaya (myininaya):

its k solve for the sqrt thing first

myininaya (myininaya):

divide both sides by 3

myininaya (myininaya):

(6x-9)^(1/2)=1

myininaya (myininaya):

to get rid of the sqrt square both sides

OpenStudy (anonymous):

I don't get it why do we divide both sides by 3

myininaya (myininaya):

We are trying to solve for x.

OpenStudy (anonymous):

okay

myininaya (myininaya):

first step to get x by itself is to divide by that 3 since that sqrt is being multiplied by that 3

myininaya (myininaya):

to undo multiplication by a number you divide by it

OpenStudy (anonymous):

okay

myininaya (myininaya):

so we have 6x-9=1

OpenStudy (anonymous):

okay

myininaya (myininaya):

+9 to both sides giving 6x=10

OpenStudy (anonymous):

okay

myininaya (myininaya):

divide 6 on both sides ( remember to undo multiplication u divide)

myininaya (myininaya):

x=10/6=5/3

myininaya (myininaya):

Don't forget to check your work

OpenStudy (anonymous):

5/3 is my answer

OpenStudy (anonymous):

how do I do that

myininaya (myininaya):

so you said x=5/3. go back to your original problem and where there is a x plug in 5/3

myininaya (myininaya):

lets see if we get the same thing on both sides and if we do that means it is a solution

myininaya (myininaya):

3*(6*5/3-9)^(1/2)=3(10-9)^(1/2)=3(1)^(1/2)=3*1=3 since the left side = the right side that means we have x=5/3 is a solution

OpenStudy (anonymous):

okay i wrote it down thank you

myininaya (myininaya):

no problem